OpenStudy (ronaldo7):

Find the radius of a circle in which a 48cm chord is 8 cm closer than a 40cm chord Does any one know how to draw and walk me through the problem I know the answer I just need help knowing how to do the problem and all the theorems that go along with it. Thank You So much for your help

OpenStudy (mathstudent55):

The 48 cm chord is 8 cm closer to what? To the center of the circle?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

 \(20^2 + (x + 8)^2 = r^2\) \(24^2 + x^2 = r^2\)